# Lightweight Joint Optimization of General-Purpose Vision-Language Models and Retrievers for RAG-Based Medical Diagnosis

### v3

@@ 4.2 Results
Table 1(a) and Table 1(b) show the effectiveness of Jomed on five medical classification and three VQA benchmarks. We compare with methods of similar computational cost; all LVLMs were not extensively pre-trained on medical data, whereas the MMed-RAG retriever was. Jomed consistently outperforms MMed-RAG, RAD, the fine-tuned RAG baseline, Jomed + PDist loss, and the Reader baseline (LVLM without retrieval).
In addition, we evaluated our model in an oracle scenario. Instead of the final fusion of logits (Section 3.3), we applied an oracle that, given the model responses for different retrieved images, selects the correct answer if it exists. [^2] For open-ended questions we do not conduct this analysis as there is no simple boundary between correct/incorrect. Our findings in Tables 1(a) and 1(b) show that the correct answer exists in a large proportion of cases, yielding superior results. This shows that in these cases, Jomed is able to surface in its top retrieved images an image which could lead to a correct prediction, however the information gets lost in the process of prediction fusion with other retrieved images. Simply taking the label with highest confidence or label with highest average confidence across retrieved images, underperforms. This motivates us to explore the use of the o3 multimodal reasoning model to detect this image. see Section 5.
@@
Jomed .91 .87 .85 .67 .62 .53 .71 .10 .64 .42 .75 .91 .75 .52
Table 5: Analysis of joint optimization effect: for inconsistent retrieval predictions, we can see that applying joint optimization boosts performance significantly for all datasets. For consistent retrieval predictions, the results are slightly improved. Note that inconsistent retrievals are challenging to all models, even non-retrieval ones.
@@
Table 6: Proportion inconsistent predictions cases, per dataset, for Jomed based on Qwen2-vl and Pixtral.
Benchmarking versus medically pretrained LVLMs for classification. In Table 3, we include previously reported results for several leading medically pretrained LVLM methods (detailed in Section 4.1). Jomed demonstrates competitive performance compared to models that underwent extensive medical pretraining (pretraining sizes summarized in Table 2). Jomed with the Pixtral backbone matches or exceeds all models except GSCo on the Breast dataset. To our knowledge, no medically pretrained LVLMs have been evaluated on the Retina benchmark, and we are not aware of any prior LVLM results on our other classification benchmarks. For binary and multi-class classification, we report only accuracy (ACC), as prior works did not report F1 scores. Conversely, for multi-label classification, we report only F1 scores. We also evaluated a medical baseline, Med-Flamingo, which generally underperformed and showed limited benefit from retrieval augmentation. We attribute this to its relatively older LLaMA backbone Touvron et al. (2023). Additionally, we experimented with using BiomedCLIP Johnson et al. (2019b) as the retriever for our general-purpose LVLMs, which exhibited a similar trend to the non-medical retriever. Full results in Appendix D.3.
Benchmarking against pretrained LVLMs for visual question answering. In Table 4, we compare our approach with medically pretrained LVLM methods (detailed in Section 4.1) for VQA. Jomed achieves competitive performance relative to extensively pretrained models. For closed-question tasks, Jomed outperforms or matches existing models. For open-ended question answering, Jomed with the Qwen2-vl backbone matches LLaVA-MedVicuna on the SLAKE dataset and LLaVA-MedLLaVA on the PathVQA dataset. Note that we do not report results on VQA-RAD, since our method was trained using an internal split of training and validation, whereas VQA-RAD provides only official training and test splits.
@@ 5 Analysis
Jomed boosts empirically challenging cases. We observed that the performance gap is substantially larger for predictions classified as inconsistent retrieval predictions (the proportion of inconsistent retrieval predictions is reported in Table 6). We define an inconsistent retrieval prediction as one in which the model predicts different labels for retrieved candidates given the same target query (see Figure 1). Specifically, if the model’s predictions vary across retrieved candidates, the instance is categorized as an inconsistent retrieval prediction. Conversely, if the model consistently predicts the same label for all retrieved candidates, it is classified as a consistent retrieval prediction.
Both the consistent and inconsistent retrieval prediction sets are established after the initial reader training stage, and their performance is assessed following joint multimodal retriever fine-tuning. Detailed results are presented in Table 5. Overall, Jomed improves performance on inconsistent retrieval predictions by +0.12 in accuracy and +0.13 in F1 score when using the Qwen2‑vl backbone, and by +0.09 in both accuracy and F1 score with the Pixtral backbone, while also offering a slight improvement on the consistent. We also observed that inconsistent retrieval predictions are empirically more challenging for the reader, FT RAG and Jomed models, with a 10–20 point performance gap between the inconsistent and consistent.
Performance with o3 as reranker. Results for inconsistent predictions remain relatively low (Table 5), even after applying joint optimization (Table 5). However, in the oracle analysis reported above, we found that Jomed can sometimes retrieve images that lead to correct predictions, together with other images that lead to wrong predictions. We thus explore whether a state-of-the-art multimodal model, when used as an optional reranker, could close the gap toward oracle performance. Specifically, we investigated whether the o3 reasoning model OpenAI (2025) could effectively select the image from the retrieved set that offers the most predictive information, guiding the model toward correct classification. We provided o3 with an image to classify, along with four retrieved images+captions, and tasked it with identifying the one containing the most predictive information.
Our results show that o3’s performance generally surpasses simply taking the image with highest confidence, but o3 is generally inferior compared to fusing logits in Jomed.We conclude that reranking in this setting remains a significant challenge even for a frontier model. Results in Appendix D.2.
Additional analyses and ablations. We evaluated the contribution of each component in our method in Appendix D.1. Additionally, we present several robustness ablations in Appendix D.4. We demonstrate robustness across different scenarios.
@@ 6 Conclusion
We demonstrated that joint optimization of general-purpose LVLMs and multimodal retrievers can achieve competitive medical diagnosis performance through lightweight fine-tuning, without medical pre-training. Future work may explore whether this approach generalizes as a cost-effective alternative to domain-specific pre-training when task-specific data and domain knowledge bases are available but large-scale pre-training is prohibitive. In addition, we identified and substantially mitigated inconsistent retrieval predictions, where different retrieved candidates lead to conflicting diagnoses. Our oracle analysis revealed a considerable performance gap between actual results and what was theoretically achievable using the retrieved images, providing a foundation for future research on closing this gap. Future work may also look into the prevalence and impact of inconsistent retrieval predictions more broadly.
@@
Sequential rather than end-to-end optimization.
We acknowledge that, in our method, gradients do not flow simultaneously between the retriever and reader modules. Nevertheless, the reader leverages signals from the retriever during training to improve its performance, while the retriever, in turn, benefits from feedback provided by the reader to refine its retrieved content. This bidirectional interaction effectively induces a joint optimization effect within our framework.
Scope of evaluation.
Our evaluation focuses on classification and visual question-answering tasks. We did not evaluate our method on report generation, which requires different capabilities (longer-form generation, clinical writing conventions) and would benefit from dedicated investigation. Additionally, while our method shows consistent improvements across diverse imaging modalities (ultrasound, fundus photography, X-ray, histopathology, dermatoscopy, CT) and tasks (visual question answering, malignant lesion classification, diabetic retinopathy severity grading), evaluation on additional specialized modalities (e.g., PET, Microscopy, OCT) and tasks (e.g., organ classification, retinal OCT disease classification) would further validate generalizability.
@@ B.2 Reader Fine-Tuning
B.3 Multimodal Retriever Fine-Tuning
@@ B.4 Evaluation Metrics
Model
Breast Derma Retina Vindr BREST VQARAD SLAKE PathVQA Mean
@@
JomedPixtral .91 .87 .80 .59 .59 .49 .56 .15 .50 .39 .77 .89 .89 .74 .50
Table 10: Performance comparison of Jomed against FT RAG baseline across medical imaging classification and VQA tasks. Both Qwen2-VL and Pixtral backbones show consistent improvements with Jomed, particularly in F1 scores. Results demonstrate that BiomedCLIP-based retrieval enhances performance similarly to the general-purpose retriever. VQA results are reported for the closed-question subset using exact match accuracy.
Model
Breast Derma Retina VQARAD
@@
Reader
only
.82 .77 .67 .54 .57 .40 .65
Jomed .81 .77 .68 .54 .60 .51 .69
Table 11: Med-Flamingo comparison with and without Jomed retrieval augmentation across medical imaging tasks. Jomed achieves best performance on 4 out of 7 metrics, with notable improvements over Reader Only in Retina classification (ACC: 0.57→0.60, F1: 0.40→0.51) and VQA-RAD (0.65→0.69). However, the overall performance gains are more modest than those observed with other backbone architectures, indicating that Med-Flamingo may have limited capacity to leverage additional retrieval context.
@@ Appendix D Additional Analysis
D.1 Analysis of model components
We analyze the contribution of each part in the model in Table 7(a) and in Table 7(b). The result of the first stage of training - the reader retrieval augmentation fine-tuning is the FT RAG then we check the contribution of training only the text retriever head and then finally is the contribution of training both the text and image retriever head as Jomed. We also explore the utility of converting the open question into a closed one in our retrieval loss. If we don’t apply o3 model to convert the open question, we only trained on the closed one, which is referred to in Table 7(b) as Closed Question Only.
D.2 Performance with a State-of-the-Art LVLM Reranker
We evaluate whether a state-of-the-art large vision–language model (LVLM), used as an optional reranker, can improve performance on inconsistent-retrieval predictions. Specifically, we test the o3 reasoning model OpenAI (2025) as a reranker: given a query image to classify and four retrieved image–caption pairs, o3 is asked to select the single pair that is most informative for predicting the correct label. We compare this reranking to four alternatives: (i) using the top-1 retrieved candidate (no reranking), (ii) selecting the model’s prediction with the highest confidence (maximum logit), (iii) using o3 with captions only (no images), and (iv) the Jomed aggregation strategy.
Our results show that o3 generally outperforms the maximum-logit baseline but remains inferior to Jomed’s logit-fusion aggregation. We evaluated performance on the Breast, Derma, and Retina datasets. For the remaining datasets—VinDr-PCXR, BRSET, and the visual question answering (VQA) datasets—we did not run this evaluation because the images and/or retrieved candidates come from restricted-access sources (e.g., MIMIC-CXR, VinDr-PCXR, and BRSET). See results in Table 9.
D.3 Jomed variants with medical pre-trained backbones
@@
D.4 Retrieval Robustness Analyses.
@@
We also test the ability of the model to still function without any retrieved context. We do this to check whether some intrinsic capability is still maintained in the model’s reader without any retrieved information. This could also be helpful in cases where model users would like it to be versatile and support both retrieval-augmented prediction and non-retrieval-augmented prediction (e.g., in cases where no retrieved contexts were discovered). As shown in Table 8 in Appendix D, removing retrieval from Jomed resulted in performance comparable or slightly lower than the reader-only model.
@@ Appendix E Prompt Design
We detail the prompts used in our experiments and analyses. To convert open questions into closed ones with the o3 model, we use the prompt in Prompt 1. To assess the utility of o3 as a re-ranker of the most predictive image–caption pair (Section D.2, Appendix D), we provide both LVLM backbones (Pixtral and Qwen2-VL) with the task-specific prompt in Prompt 2.
We also include the training prompts for Jomed across datasets—Prompts 3, 5, 4, 7, 6—and the VQA prompt (Prompt 8).

### v4

@@ 4.2 Results
Table 1(a) and Table 1(b) show the effectiveness of CLARE on five medical classification and three VQA benchmarks. We compare with methods of similar computational cost; all LVLMs were not extensively pre-trained on medical data, whereas the MMed-RAG retriever was. CLARE consistently outperforms MMed-RAG, RAD, the fine-tuned RAG baseline, CLARE + PDist loss, and the Reader baseline (LVLM without retrieval). In addition, we evaluated our model in an oracle scenario. Instead of the final fusion of logits (Section 3.3), we applied an oracle that, given the model responses for different retrieved images, selects the correct answer if it exists. [^2] For open-ended questions we do not conduct this analysis as there is no simple boundary between correct/incorrect. Our findings in Tables 1(a) and 1(b) show that the correct answer exists in a large proportion of cases, yielding superior results. This shows that in these cases, CLARE is able to surface in its top retrieved images an image which could lead to a correct prediction, however the information gets lost in the process of prediction fusion with other retrieved images. Simply taking the label with highest confidence or label with highest average confidence, underperforms. This motivates us to explore the o3 multimodal reasoning model to detect this image (Section 5).
@@
CLARE .91 .87 .85 .67 .62 .53 .71 .10 .64 .42 .75 .91 .75 .52
Table 5: Analysis of LVLM-aware multimodal retrieval fine-tuning effect: for inconsistent retrieval predictions, we can see that applying LVLM-aware multimodal retrieval fine-tuning boosts performance significantly for all datasets. For consistent retrieval predictions, the results are slightly improved. Note that inconsistent retrievals are challenging to all models, even non-retrieval ones.
@@
Table 6: Proportion inconsistent predictions cases, per dataset, for CLARE based on Qwen2-vl and Pixtral.
Benchmarking versus medically pretrained LVLMs for classification. In Table 3, we include previously reported results for leading medically pretrained LVLM methods (detailed in Section 4.1). CLARE demonstrates results competitive with models that underwent extensive medical pretraining (pretraining sizes in Table 2). CLARE with the Pixtral backbone matches or exceeds all models except GSCo on the Breast dataset. To our knowledge, no medically pretrained LVLMs have been evaluated on the Retina benchmark, and we are not aware of any prior LVLM results on our other classification benchmarks. For binary and multi-class classification, we report only accuracy (ACC), as prior works did not report F1 scores. Conversely, for multi-label classification, we report only F1 scores. We also evaluated a medical baseline, Med-Flamingo, which generally underperformed and showed limited benefit from retrieval augmentation. We attribute this to its relatively older LLaMA backbone Touvron et al. (2023). Additionally, we experimented with using BiomedCLIP Johnson et al. (2019b) as the retriever for our general-purpose LVLMs, which exhibited a similar trend. Full results in Appendix D.4.
Benchmarking against pretrained LVLMs for visual question answering. In Table 4, we compare our approach with medically pretrained LVLM methods (detailed in Section 4.1) for VQA. CLARE achieves competitive performance relative to extensively pretrained models. For closed-question tasks, CLARE outperforms or matches existing models. For open-ended question answering, CLARE with the Qwen2-vl backbone matches LLaVA-MedVicuna on the SLAKE dataset and LLaVA-MedLLaVA on the PathVQA dataset. Note that we do not report results on VQA-RAD, since our method was trained using an internal split of training and validation, whereas VQA-RAD provides only official training and test splits.
@@ 5 Analysis
CLARE boosts empirically challenging cases. We observe that the performance gap is substantially larger for predictions classified as inconsistent retrieval predictions (Table 6 shows the proportion of such cases). We define an inconsistent retrieval prediction as one in which the model predicts different labels for retrieved candidates given the same query (see Figure 1), i.e., if the model’s predictions vary across retrieved candidates. Conversely, if the model predicts the same label for all retrieved candidates, the instance is consistent.
Consistent and inconsistent prediction sets are established after the initial reader training stage, and their performance is assessed following LVLM-aware multimodal retrieval fine-tuning. Detailed results are presented in Table 5. Overall, CLARE improves performance on inconsistent retrieval predictions by +0.12 in accuracy and +0.13 in F1 score when using the Qwen2‑vl backbone, and by +0.09 in both accuracy and F1 score with the Pixtral backbone, while also offering a slight improvement on the consistent. We also observed that inconsistent retrieval predictions are empirically more challenging for the reader, FT RAG and CLARE models, with a 10–20 point performance gap.
Performance with reranking. Results for inconsistent predictions remain relatively low (Table 5). However, in the oracle analysis reported above, we found that CLARE can sometimes retrieve images that lead to correct predictions, together with other images that lead to wrong predictions. We thus explore whether a state-of-the-art multimodal model, when used as an optional reranker, could close the gap toward oracle performance. Specifically, we investigated whether the o3 reasoning model OpenAI (2025) could select a retrieved image that leads to correct prediction. We provided o3 with an image to classify, along with four retrieved images+captions, and tasked it with identifying the one containing the most predictive information. Our results show that o3’s performance generally surpasses simply taking the image with highest confidence, but o3 is generally inferior compared to fusing logits in CLARE. We conclude that reranking in this setting remains a significant challenge even for a frontier model. Results in Appendix D.3.
Retrieval relevance and additional analyses. We explore the utility of our method in retrieving more relevant candidates. In our setting where no retrieval ground-truth labels are available, we use GPT-5.2 as a judge to compare relevance of retrieved candidates before and after the LVLM-aware retriever fine-tuning. Our method improves candidate relevance with an average 17% win rate, as detailed in Appendix D.1. We further evaluate the contribution of each component in Appendix D.2 and present robustness ablations in Appendix D.5.
@@ 6 Conclusion
We demonstrated that LVLM-aware multimodal retrieval achieves competitive medical diagnosis performance through lightweight, data-efficient fine-tuning, without medical pre-training. Future work may explore whether this approach generalizes as a cost-effective alternative to domain-specific pre-training when task-specific data and domain knowledge bases are available but large-scale pre-training is prohibitive. In addition, we identified and substantially mitigated inconsistent retrieval predictions, where different retrieved candidates lead to conflicting diagnoses. Our oracle analysis revealed a considerable performance gap between actual results and what was theoretically achievable using the retrieved images, providing a foundation for future research on closing this gap. Future work may also look into the prevalence and impact of inconsistent retrieval predictions more broadly.
@@
Scope of evaluation. Our evaluation focuses on classification and visual question-answering tasks, including both open questions (a text-generation setting) and closed questions (closed-form answer selection). We did not evaluate our method on report generation, which requires different capabilities (longer-form generation, clinical writing conventions) and would benefit from dedicated investigation. Additionally, while our method shows consistent improvements across diverse imaging modalities (ultrasound, fundus photography, X-ray, histopathology, dermatoscopy, CT) and tasks (visual question answering, malignant lesion classification, diabetic retinopathy severity grading), evaluation on additional specialized modalities (e.g., PET, Microscopy, OCT) and tasks (e.g., organ classification, retinal OCT disease classification) would further validate generalizability.
@@ B.2 Reader Fine-Tuning
B.3 LVLM-Aware Multimodal Retrieval Fine-Tuning
@@ B.4 Evaluation Metrics
Model Breast Derma Retina Vindr BREST VQARAD SLAKE PathVQA Mean
@@
CLAREPixtral .91 .87 .80 .59 .59 .49 .56 .15 .50 .39 .77 .89 .89 .74 .50
Table 10: Performance comparison of CLARE against FT RAG baseline across medical imaging classification and VQA tasks. Both Qwen2-VL and Pixtral backbones show consistent improvements with CLARE, particularly in F1 scores. Results demonstrate that BiomedCLIP-based retrieval enhances performance similarly to the general-purpose retriever. VQA results are reported for the closed-question subset using exact match accuracy.
Model Breast Derma Retina VQARAD
@@
Reader only .82 .77 .67 .54 .57 .40 .65
CLARE .81 .77 .68 .54 .60 .51 .69
Table 11: Med-Flamingo comparison with and without CLARE retrieval augmentation across medical imaging tasks. CLARE achieves best performance on 4 out of 7 metrics, with notable improvements over Reader Only in Retina classification (ACC: 0.57→0.60, F1: 0.40→0.51) and VQA-RAD (0.65→0.69). However, the overall performance gains are more modest than those observed with other backbone architectures, indicating that Med-Flamingo may have limited capacity to leverage additional retrieval context.
@@ C.2 Medical Pre - trained baseline models
Backbone Winner Breast Derma Retina VinDr BRSET VQA-RAD SLAKE PathVQA Mean
Qwen2-vl Tie .20 .59 .23 .25 .50 .24 .50 .20 .34
FT RAG Wins .30 .09 .32 .36 .24 .25 .16 .35 .26
CLARE Wins .50 .30 .44 .46 .25 .50 .34 .45 .41
Pixtral Tie .24 .51 .51 .47 .35 .12 .15 .29 .33
FT RAG Wins .26 .12 .15 .19 .40 .32 .37 .30 .26
CLARE Wins .50 .35 .34 .34 .52 .54 .47 .41 .43
Table 12: Analysis of relevance of retrieval candidates before and after LVLM-aware multimodal retrieval fine-tuning. Training the multimodal retrieval significantly improves performance, with a mean improvement of 0.17 across Qwen2-vl and Pixtral.
@@ Appendix D Additional Analysis
D.1 Retrieval Relevancy Analysis
We explore whether LVLM-aware multimodal retrieval fine-tuning improves the quality of retrieval candidates compared to those produced before fine-tuning. Importantly, our setup does not include ground-truth retrieval labels; instead, the retriever learns which candidates better steer the frozen LVLM toward the correct downstream prediction. To measure candidate improvement, we compare the retriever’s image and text heads before vs. after fine-tuning in a head-to-head ranking task. We then prompt GPT-5.2 to judge which candidate more closely resembles the ground-truth prediction, assuming that higher similarity to the ground-truth prediction indicates better evidence steering. When both candidates are identical, we record a tie. To avoid any location biases, we conduct shuffling between entering the model and reducing cost; we use a subset of 10% of the samples. Prompt 2 details the prompt for the task. Table 12 shows that LVLM-aware multimodal retrieval fine-tuning improves candidate relevance, with a mean gain of 0.17 for both Qwen2-VL and Pixtral, demonstrating the practical utility of our method.
D.2 Analysis of Model Components
We analyze the contribution of each part in the model in Table 7(a) and in Table 7(b). The result of the first stage of training - the reader retrieval augmentation fine-tuning is the FT RAG then we check the contribution of training only the text retriever head and then finally is the contribution of training both the text and image retriever head as CLARE. We also explore the utility of converting the open question into a closed one in our retrieval loss. If we don’t apply o3 model to convert the open question, we only trained on the closed one, which is referred to in Table 7(b) as Closed Question Only.
D.3 Performance with a State-of-the-Art LVLM Reranker
We evaluate whether a state-of-the-art large vision–language model (LVLM), used as an optional reranker, can improve performance on inconsistent-retrieval predictions. Specifically, we test the o3 reasoning model OpenAI (2025) as a reranker: given a query image to classify and four retrieved image–caption pairs, o3 is asked to select the single pair that is most informative for predicting the correct label. We compare this reranking to four alternatives: (i) using the top-1 retrieved candidate (no reranking), (ii) selecting the model’s prediction with the highest confidence (maximum logit), (iii) using o3 with captions only (no images), and (iv) the CLARE aggregation strategy.
Our results show that o3 generally outperforms the maximum-logit baseline but remains inferior to CLARE’s logit-fusion aggregation. We evaluated performance on the Breast, Derma, and Retina datasets. For the remaining datasets—VinDr-PCXR, BRSET, and the visual question answering (VQA) datasets—we did not run this evaluation because the images and/or retrieved candidates come from restricted-access sources (e.g., MIMIC-CXR, VinDr-PCXR, and BRSET). See results in Table 9.
Backbone Model Breast Derma Retina VinDr BRSET
ACC F1 ACC F1 ACC F1 ACC F1 ACC F1
Qwen2-vl CLARE N=2 .87 .83 .74 .50 .63 .50 .56 .14 .47 .35
CLARE N=4 .87 .84 .76 .50 .65 .50 .57 .14 .49 .37
CLARE N=6 .87 .84 .75 .50 .63 .52 .58 .14 .48 .37
Pixtral CLARE N=2 .89 .86 .80 .63 .58 .48 .54 .13 .49 .35
CLARE N=4 .90 .87 .80 .62 .60 .51 .56 .14 .51 .37
CLARE N=6 .90 .87 .80 .63 .61 .51 .55 .13 .49 .36
Table 13: Ablation on the number of retrieved candidates (N) for classification tasks.
Backbone Model VQA-RAD SLAKE PathVQA
Closed Open Closed Open Closed Open
Qwen2-vl CLARE N=2 .78 .46 .88 .82 .92 .37
CLARE N=4 .79 .48 .90 .84 .93 .38
CLARE N=6 .78 .47 .89 .83 .93 .38
Pixtral CLARE N=2 .78 .45 .89 .83 .93 .38
CLARE N=4 .78 .47 .90 .84 .93 .37
CLARE N=6 .78 .47 .89 .84 .93 .38
Table 14: Ablation on the number of retrieved candidates (N) for visual question answering tasks.
D.4 CLARE Variants with Medical Pre-Trained Backbones
@@
D.5 Retrieval Robustness Analyses.
@@
We also test the ability of the model to still function without any retrieved context. We do this to check whether some intrinsic capability is still maintained in the model’s reader without any retrieved information. This could also be helpful in cases where model users would like it to be versatile and support both retrieval-augmented prediction and non-retrieval-augmented prediction (e.g., in cases where no retrieved contexts were discovered). As shown in Table 8 in Appendix D, removing retrieval from CLARE resulted in performance comparable or slightly lower than the reader-only model.
@@
D.6 Different number of retrieved candidates
Our model demonstrates robustness across varying numbers of retrieved candidates. We explore performance for N = 2, 4, and 6, showing consistent results across the different configurations. Results are presented in Table 13 for classification tasks and in Table 14 for visual question answering tasks.
@@ Appendix E Prompt Design
We detail the prompts used in our experiments and analyses. To convert open questions into closed ones with the o3 model, we use the prompt in Prompt 1. To assess the utility of o3 as a re-ranker of the most predictive image–caption pair (Section D.3, Appendix D), we provide both LVLM backbones (Pixtral and Qwen2-VL) with the task-specific prompt in Prompt 3.
We also include the training prompts for CLARE across datasets—Prompts 4, 6, 5, 8, 7—and the VQA prompt (Prompt 9).
@@
You are an expert medical AI assistant specializing in evaluating medical image and text relevance for clinical decision-making.
Compare two medical (image, text) pairs and determine which is MORE RELEVANT to the ground truth label.
Ground Truth Label: "{gt_text}"
First Image: {image_1}
First Text: {text_1}
Second Image: {image_2}
Second Text: {text_2}
Consider:
- Medical image findings
- Medical terminology accuracy in text
- Clinical relevance to the ground truth
- Diagnostic information alignment
- How useful each (image, text) pair would be for the given condition
Return JSON format:
"choice": <1 or 2>,
"confidence": <"high", "medium", or "low">,
"explanation": "<brief 1-sentence explanation>"
Be concise and accurate. Choose the option that is MORE medically relevant to the ground truth.
Listing 2: Prompt used for retrieval candidate relevance evaluation